# MR-CLIP: Efficient Metadata-Guided Learning of MRI Contrast Representations

### v1

@@ 1 Introduction
Magnetic resonance imaging (MRI) is the cornerstone of modern neuroimaging, offering diverse contrast mechanisms sensitive to different tissue properties and pathologies. However, in clinical settings, the interpretation and organization of MR images is challenged by substantial variability in acquisition protocols, scanner vendors, and data completeness [15]. Unlike curated research datasets that are relatively homogeneous, real-world clinical data is highly variable and often lack explicit sequence labels such as T1-weighted or T2-weighted [3]. Even when present, these labels offer only a coarse approximation of image contrast. Instead, the true visual characteristics of an MR image are determined by acquisition settings, such as scanner type, echo time, and repetition time, which are embedded within the Digital Imaging and Communications in Medicine (DICOM) [11] metadata. Reliable and complete metadata are essential for clinical interpretation, image retrieval, and picture routing and communication systems (PACS) routing and archiving. Furthermore, robust contrast-aware representations are foundational for clinical applications such as cross-site and cross-modality data harmonization, and modality-invariant representations [20, 13].
Recent studies highlight the importance of metadata in shaping MR image appearance; for example, ContextMRI [1] incorporates DICOM metadata into image reconstruction, showing that contrast-aware models can improve image quality. Similarly, a contrastive learning approach for retinal OCT images uses patient metadata to enhance representations [4]. TUMSyn [17] synthesizes MR images guided by DICOM tags using contrastive learning; however, it relies primarily on controlled research datasets with limited scanner and protocol diversity. Other works [10, 2, 3, 9] use supervised learning to estimate metadata or sequence type from images or tags, though they mainly predict high-level labels such as T1/T2 contrast while overlooking fine-grained MR sequences parameter variability. CLIP [14] learns aligned image–text representations via contrastive learning, enabling powerful zero-shot transfer across diverse tasks. In medicine, models such as BiomedCLIP [19], PMC-CLIP [18], and MedCLIP [16] adapt this framework to medical data for retrieval and classification. However, they primarily focus on general tasks, overlooking domain-specific MR acquisition metadata that critically shapes image contrast and appearance.
@@ 2 Methods
We construct a structured label space derived from key DICOM metadata fields influencing MR image contrast supervise to align images and metadata without relying on manual labels. These include categorical tags, including Manufacturer, Scanner Model, Imaging Plane, Field Strength, Sequence Type, Sequence Variant, Flip Angle, and numerical tags Echo Time (TE), Repetition Time (TR), and Inversion Time (TI). The distributions of these metadata fields are illustrated in Fig. 2. Each contrast-aware label corresponds to a unique combination of grouped metadata values. Categorical tags (Fig. 2a) are grouped by exact values. Numerical tags are discretised to reduce the sensitivity to minor acquisition differences that do not meaningfully affect the image appearance (Fig. 2b). TE and TR are jointly quantized into a $20\times 20$ grid to reflect clinically relevant ranges, with TI binned separately as >85% scans lack inversion pulses and the remainder follow a sparse but structured distribution. While the Series Description field can offer useful context, it is excluded from label construction due to site-dependent variability; instead, it is used only as an auxiliary input.
Unlike traditional CLIP, which uses one-to-one image-text pairs with an InfoNCE loss [12], MR-CLIP employs a supervised contrastive (SupCon) loss [7] that supports many-to-many positives. This is well-suited to MRI, where small variations in acquisition parameters (e.g., TE, TR) often do not affect the underlying contrast. While CLIP would penalize these near-identical cases due to strict pairings, SupCon treats all samples sharing the same grouped contrast label as positives, enabling more flexible and semantically consistent supervision. Moreover, by aligning slices from different subjects and anatomical locations, MR-CLIP aims to learn contrast representations that are invariant to anatomy and subject-specific variation. Formally, given a batch of embeddings ${z_{i}}_{i=1}^{N}$, the SupCon loss for an anchor embedding $z_{i}$ is defined as:
@@
Implementation Details: MR-CLIP is implemented in PyTorch and trained on 3 NVIDIA A100 GPUs (40GB), using a batch size of 3000 for each GPU with sharding loss following CLIP implementation [6]. We optimize with Adam ($\text{lr}=1\mathrm{e}{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.98$) and a weight decay of 0.2, training for 100 epochs with 2000 warm-up steps. To reduce memory usage and scale efficiently, we enable gradient checkpointing, and local loss computation. Patch dropout (0.5) and text dropout (0.2) are applied, along with standard image augmentations (random affine, resized crop, Gaussian blur, horizontal flip).
@@ 3 Experiments
Cross-modal Retrieval and Linear Classification: We investigate the impact of loss functions and architectural choices on representation quality, and additionally compare MR-CLIP to several BiomedCLIP variants, a model pre-trained on large-scale biomedical image-text pairs, as summarized in Table 1. We consider three BiomedCLIP configurations: (1) the original frozen model, (2) a fully fine-tuned version, and (3) a variant with only the final layer trained. We evaluate models on two tasks: (i) cross-modal retrieval across three settings: retrieving metadata from 2D slices (image-to-text), from 3D scans via majority voting across slices (3D scan-to-text), and retrieving images from metadata (text-to-image); and (ii) linear probe classification, where a simple linear classifier is trained on frozen image embeddings to predict metadata labels. MR-CLIP achieves the highest (R@1) in both 3D scan-to-text (78.7%) and text-to-image (90.9%) retrieval, significantly outperforming the best BiomedCLIP variant (67.4% R@1 for 3D scan-to-text). While InfoNCE achieves higher image-to-text R@5 and R@10 scores, likely due to its one-to-one training supervision, MR-CLIP excels overall. Furthermore, MR-CLIP attains the highest linear classification accuracy (82.6%), indicating that its learned representations are more discriminative and better aligned with metadata.
@@
BiomedCLIP - Full 50.0 78.5 82.6 67.4 89.1 92.1 38.5 65.8 71.8 75.5
BiomedCLIP - Last Layer 53.2 78.6 83.1 68.3 90.0 92.7 43.3 68.1 72.9 75.6

### v2

@@ 1 Introduction
Magnetic resonance imaging (MRI) is the cornerstone of modern neuroimaging, offering diverse contrast mechanisms sensitive to different tissue properties and pathologies. However, in clinical settings, the interpretation and organization of MR images is challenged by substantial variability in acquisition protocols, scanner vendors, and data completeness [15]. Unlike curated research datasets that are relatively homogeneous, real-world clinical data is highly variable and often lack explicit sequence labels such as T1-weighted or T2-weighted [3]. Even when present, these labels offer only a coarse approximation of image contrast. Instead, the true visual characteristics of an MR image are determined by acquisition settings, such as scanner type, echo time, and repetition time, which are embedded within the Digital Imaging and Communications in Medicine (DICOM) [11] metadata. Reliable and complete metadata are essential for clinical interpretation, image retrieval, and picture archiving and communication systems (PACS) routing and archiving. Furthermore, robust contrast-aware representations are foundational for clinical applications such as cross-site and cross-modality data harmonization, and modality-invariant representations [20, 13].
Recent studies highlight the importance of metadata in shaping MR image appearance; for example, [1] incorporates DICOM metadata into image reconstruction, showing that contrast-aware models can improve image quality. Similarly, a contrastive learning approach for retinal OCT images uses patient metadata to enhance representations [4]. TUMSyn [17] synthesizes MR images guided by DICOM tags using contrastive learning; however, it relies primarily on controlled research datasets with limited scanner and protocol diversity. Other works [10, 2, 3, 9] use supervised learning to estimate metadata or sequence type from images or tags, though they mainly predict high-level labels such as T1/T2 contrast while overlooking fine-grained MR sequences parameter variability. CLIP [14] learns aligned image–text representations via contrastive learning, enabling powerful zero-shot transfer across diverse tasks. In medicine, models such as BiomedCLIP [19], PMC-CLIP [18], and MedCLIP [16] adapt this framework to medical data for retrieval and classification. However, they primarily focus on general tasks, overlooking domain-specific MR acquisition metadata that critically shapes image contrast and appearance.
@@ 2 Methods
To guide alignment between MR images and metadata without manual labeling, a structured label space is constructed using critical DICOM metadata fields that jointly influence image contrast. These fields include categorical tags, including Manufacturer, Scanner Model, Imaging Plane, Field Strength, Sequence Type, Sequence Variant, Series Description, Flip Angle, and numerical tags Echo Time (TE), Repetition Time (TR), and Inversion Time (TI). Fig. 2 illustrates the distributions of the metadata fields. Each contrast-aware label corresponds to a unique combination of these metadata values. Categorical tags (Fig. 2a) are grouped by exact values. Numerical tags are discretised to reduce the sensitivity to minor acquisition differences that do not meaningfully affect the image appearance (Fig. 2b). TE and TR are jointly quantized into a $\times 20$ 20 grid covering typical ranges observed in clinical protocols, with TI binned separately as >85% scans lack inversion pulses and the remainder follow a sparse but structured distribution. While the Series Description field can offer useful context, it is excluded from label construction due to site-dependent variability; instead, it is used only as an auxiliary input.
Unlike traditional CLIP, which uses one-to-one image-text pairs with an InfoNCE loss [12], MR-CLIP employs supervised contrastive (SupCon) loss [7] that supports many-to-many positives. This is well-suited to MRI, where small variations in acquisition parameters (e.g., TE, TR) do not affect the underlying contrast. While CLIP would penalize these near-identical cases due to strict pairings, SupCon treats all samples sharing the same grouped contrast label as positives, enabling more flexible and semantically consistent supervision. Moreover, by aligning slices from different subjects and anatomical locations, MR-CLIP aims to learn contrast representations that are invariant to anatomy and subject-specific variation. Formally, given a batch of embeddings ${z_{i}}_{i=1}^{N}$, the SupCon loss for an anchor embedding $z_{i}$ is defined as:
@@
Implementation Details: MR-CLIP is implemented in PyTorch and trained on 3 NVIDIA A100 GPUs (40GB), using a batch size of 3000 for each GPU with sharding loss following CLIP implementation [6]. We optimize with Adam ($\text{lr}=1\mathrm{e}{-4}$, $\beta_{1}=0.9$, $\beta_{2}=0.98$) and a weight decay of 0.2, training for 100 epochs with 2000 warm-up steps. Gradient checkpointing is enabled to reduce memory usage. Patch dropout (0.5) and text dropout (0.2) are applied, along with standard image augmentations (random affine, resized crop, Gaussian blur, horizontal flip).
@@ 3 Experiments
Cross-modal Retrieval and Linear Classification: We investigate the impact of loss functions and architectural choices on representation quality, and additionally compare MR-CLIP to BiomedCLIP, a vision–language model pre-trained on large-scale biomedical image–text pairs, in both its original frozen and fully fine-tuned configurations, as summarized in Table 1. We evaluate models on two tasks: (i) cross-modal retrieval across three settings: retrieving metadata from 2D slices (image-to-text), from 3D scans via majority voting across slices (3D scan-to-text), and retrieving images from metadata (text-to-image); and (ii) linear probe classification, where a simple linear classifier is trained on frozen image embeddings to predict metadata labels. MR-CLIP achieves the highest (R@1) in both 3D scan-to-text (78.7%) and text-to-image (90.9%) retrieval, significantly outperforming the the fine-tuned BiomedCLIP (67.4% R@1 for 3D scan-to-text). While InfoNCE achieves higher image-to-text R@5 and R@10 scores, likely due to its one-to-one training supervision, MR-CLIP demonstrates the strongest overall performance. Furthermore, MR-CLIP attains the highest linear classification accuracy (82.6%), indicating that its learned representations are more discriminative and better aligned with metadata.
@@
BiomedCLIP (Fine-tuned) 50.0 78.5 82.6 67.4 89.1 92.1 38.5 65.8 71.8 75.5